### v2

# HI-SQL: Optimizing Text-to-SQL Systems through Dynamic Hint Integration

Abstract
Text-to-SQL generation bridges the gap between natural language and databases, enabling users to query data without requiring SQL expertise. While large language models (LLMs) have significantly advanced the field, challenges remain in handling complex queries that involve multi-table joins, nested conditions, and intricate operations. Existing methods often rely on multi-step pipelines that incur high computational costs, increase latency, and are prone to error propagation. To address these limitations, we propose HI-SQL, a pipeline that incorporates a novel hint generation mechanism utilizing historical query logs to guide SQL generation. By analyzing prior queries, our method generates contextual hints that focus on handling the complexities of multi-table and nested operations. These hints are seamlessly integrated into the SQL generation process, eliminating the need for costly multi-step approaches and reducing reliance on human-crafted prompts. Experimental evaluations on multiple benchmark datasets demonstrate that our approach significantly improves query accuracy of LLM-generated queries while ensuring efficiency in terms of LLM calls and latency, offering a robust and practical solution for enhancing Text-to-SQL systems.
I Introduction
The task of converting natural language queries into SQL statements, known as Text-to-SQL parsing, has consistently been a major area of interest within both the natural language processing and database communities. This technology enables users to interact with databases without requiring proficiency in SQL, thereby democratizing data access and facilitating informed decision-making across various domains [b9].
Early Text-to-SQL systems relied heavily on rule-based methods and human-engineered features, which, while effective to a degree, often struggled with scalability and adaptability to diverse database schemas [b10]. The advent of deep learning and neural network models marked a significant shift, introducing sequence-to-sequence architectures that improved the translation of natural language to SQL [b11].
Current advancements in large language models (LLMs) have further propelled the field, enhancing the understanding of natural language semantics and the generation of accurate SQL queries. Recent studies have highlighted the impact of LLMs on Text-to-SQL parsing, noting their ability to capture complex linguistic patterns and improve performance across benchmarks [b12].
The development of large-scale human-labeled data sets, such as Spider [b13] and BIRD [b14], has been instrumental in evaluating and advancing Text-to-SQL systems. These datasets provide diverse and complex queries that test the robustness and adaptability of current models.
Despite significant advancements, generating accurate SQL queries for complex operations that span multiple tables, nested conditions, and intricate joins continues to be a challenge within Text-to-SQL systems. Recent efforts have focused on leveraging LLMs to address these issues by designing agentic systems that combine multi-step processes with verification methods for generating SQL queries [b8, b15, b16]. While these approaches have demonstrated effectiveness in improving query accuracy, they also lead to increased computational overhead due to the necessity of chaining multiple LLM calls. This results in high token usage and latency, while also making the system vulnerable to error propagation, as inaccuracies in one stage can negatively impact subsequent steps, ultimately compromising the system’s reliability.
In this paper, we propose HI-SQL – a system that leverages historical query logs to enhance the Text-to-SQL translation process. By systematically analyzing previously executed queries, our method generates contextual hints designed to address the complexities of handling multi-table queries, nested operations, and intricate conditions. Then, these hints are seamlessly integrated into the SQL generation pipeline, guiding the system toward producing more accurate and efficient queries. This targeted approach directly addresses one of the primary weaknesses of LLM-generated queries — handling complex database operations — while simplifying the overall system architecture and improving efficiency.
Unlike traditional agentic systems, such as CHASE-SQL [b15], which often require labor-intensive human annotations or expert-crafted few-shot examples to design effective prompts, our approach automates the hint generation process. By eliminating the need for human or domain-specific input in prompt design, our method significantly reduces manual effort and enhances scalability and adaptability across various database environments.
We rigorously evaluate our approach across multiple benchmarks, including SPIDER, BIRD, and ACME Insurance[b29], to assess its effectiveness in diverse scenarios. Experimental results reveal substantial improvements over baseline models, and results comparable to open-source state-of-the-art methods in the respective benchmarks. demonstrating the value of incorporating query-based hints for enhancing the performance of Text-to-SQL systems. Our solution achieves better query accuracy while reducing computational expenses by minimizing the number of LLM calls required, making it an efficient alternative to existing state-of-the-art methods. The main contributions of this work are two-fold viz: (i). We introduce an automatic hint generation mechanism using historical query logs to improve SQL generation for complex queries; and (ii). We present a SQL generation pipeline that reduces computational cost and latency by minimizing LLM calls while maintaining accuracy comparable to the state-of-the-art methods.
These contributions underscore the importance of leveraging historical data insights for developing more efficient, accurate, and scalable Text-to-SQL systems, offering promising directions for future research in natural language interfaces for databases.
Figure 1: Hint Curation Pipeline: The hint curation process utilizes historical queries and SQL Verifier to generate hints.
Figure 2: SQL Generation Pipeline: This pipeline facilitates the SQL query generation process using previously generated hints and the SQL Verification process using the SQL Verifier pipeline.
II Related Works
Text-to-SQL generation has evolved significantly, transitioning from early sequence-to-sequence models to advanced transformer-based architectures. Initial methods employed Recurrent Neural Networks (RNNs) and Long Short-Term Memory (LSTM) networks for encoding user queries and database schemas, utilizing slot-filling or auto-regressive decoding techniques to construct SQL queries. For instance, Seq2SQL [b17] introduced reinforcement learning to enhance query generation, while SQLNet [b18] employed a sketch-based approach to predict query structures before filling in details, thereby improving accuracy and efficiency. TypeSQL [typesql] extended SQLNet by incorporating type information, enabling better handling of complex queries involving varied data types.
Advancements in the field of Text-to-SQL have led to the development of models like IRNet [b19], which introduced a graph-based encoder to capture relationships between database schema elements and user queries, improving performance on complex queries. RAT-SQL [b20] further refined this approach by using a relation-aware transformer for schema encoding to manage multi-table queries. Meanwhile, SQLova [b21] achieved significant progress by integrating BERT-style models to enhance schema representation and column attention, setting new benchmarks for Text-to-SQL systems.
Transformer-based models have played a pivotal role in advancing Text-to-SQL parsing. The T5 [b22] model family, including T5-3B, demonstrated strong performance when fine-tuned on Text-to-SQL tasks. Domain-specific adaptations such as MedT5SQL [b23] for healthcare and EDU-T5 for education showcased the potential of tailored models to address unique challenges in specialized fields. Additionally, PICARD [b24] introduced constrained auto-regressive decoding, improving performance on dialogue-based and multi-turn SQL generation tasks.
The advent of large language models (LLMs) has redefined the field, with techniques like schema linking, self-correction, and self-consistency enhancing query accuracy for complex tasks. Benchmarks such as as Spider, BIRD, and other domain-specific datasets have facilitated the evaluation of these systems. More recently, specialized SQL agents have become one of the standard techniques for generating SQL queries. These agents often consist of multiple LLM calls for steps like entity disambiguation, schema linking, multi-SQL generation, self-correction/evolution etc [b8, b26]. These agentic frameworks for SQL generation rely on multiple calls to LLMs and external tools, introducing latency and cost. Such frameworks often incorporate human-curated domain knowledge or few-shot prompts for enhanced performance. Another recent work [b27] utilizes the human-curated ontology of the schema to generate cypher queries and then convert them back to SQL queries.
Despite the current advancement, challenges remain in handling highly complex queries and improving efficiency. Our work addresses these gaps by leveraging historical query logs to generate contextual hints, enhancing accuracy, and reducing computational costs in Text-to-SQL systems.
III Methodology
Our pipeline consists of three modules, namely (i). Hint Curation, (ii). SQL Generation and (iii). SQL Verifier. In this section we describe each of these modules in detail.
III-A Hint Curation
The Hint Curation (c.f Figure 1) process leverages a large language model (LLM) to select and refine a diverse set of historical queries that reflect varying levels of complexity. These queries encompass tasks such as managing relationships between multiple tables, employing advanced data linking methods, applying conditional filtering, and performing computational operations like aggregations and averages. By relying on the LLM, this approach ensures a comprehensive exploration of diverse query structures and challenges commonly faced in database systems, minimizing the need for manual selection. Historical queries, along with their associated database schemas, are combined into structured prompts that guide the LLM in analyzing and identifying representative examples of complex query patterns. These prompts assist the LLM in extracting queries that showcase critical relationships between tables and address broader challenges in query design, including dynamic filtering, recursive structures, and integrating data from multiple sources. The extracted queries are run through the hosted database to verify their syntactical correctness. If any queries are found to be incorrect, the LLM automatically corrects them, thereby reducing reliance on manual review and correction. In addition to handling multi-table relationships and filtering, the LLM is tasked with recognizing more advanced complexities, such as hierarchical data structures, and analytical computations (e.g., row-based calculations), as well as set operations like unions, intersections, and differences. It identifies key patterns in table joins for a database schema, such as equi-joins, non-equi-joins, and self-joins, while also understanding how grouping and sorting enhance query outputs.
The final curated set of SQL queries (c.f Table III-C), accompanied by detailed descriptions, provides actionable examples that improve the LLM’s ability to handle advanced query design. By automating this hint-generation process, the need for human-driven tasks such as creating ontologies or crafting few-shot examples for multi-agent systems-is significantly reduced. This ensures a faster and more scalable approach to enhancing the LLM’s understanding of the complexities in query construction while also improving performance.
III-B SQL Generation
For SQL generation (c.f Figure 2), the LLM is provided with hints extracted through the Hint Curation process, along with the database schema and the Natural Language Query (NLQ), all formatted within a structured prompt. With the growing context window of modern LLMs, we do not utilise a query-specific schema filtering or linking module, thereby reducing the error incurred during schema linking [b28]. This enables the LLM to generate an SQL query corresponding to the NLQ. The generated SQL is then processed through an SQL verifier to ensure its correctness.
III-C SQL Verification
The SQL verification pipeline ensures the correctness of the generated SQL by executing it on a hosted database and checking for both results and errors. If errors are encountered, the error message, along with the schema and the NLQ, is fed back into the LLM to generate a corrected SQL query. To address the non-deterministic and hallucinatory tendencies of LLMs, a retry loop has been implemented. If the LLM fails to correct the query, the verification process is retried up to a maximum of $C$ attempts. If all retries fail, the pipeline logs the error along with the NLQ for further manual intervention.
TABLE I: A sample of hints for the financial database from BIRD, generated by our hint curation pipeline.